OpenStudy (anonymous):

I need some help with a word problem, I just need someone to check my answer for me, so if anyone is up to it, please and thank you! :) I will post the question and my answer below...

OpenStudy (anonymous):

The Question: Nakim simplified 3√2x+x√8x-5√18x got -10x√2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression.

OpenStudy (anonymous):

My Answer: Part 1: Nakim would have missed the x in the middle term, and had combined all to get -10 sqrt(2x) but the solution actually is 2 (sqrt(2x))(x-6) Part 2: First, you have to simplify the expression 3sqrt(2x)+ xsqrt(8x) - 5sqrt(18x) 8=2^2*2 18=3^2*2 so , the expression is now: 3sqrt(2x) + xsqrt(4*2x) - 5sqrt(9*2x) = sqrt(2x) + 2xsqrt(2x) - 15sqrt(2x) Is this right?

OpenStudy (anonymous):

With the solution being: 2 (sqrt(2x))(x-6)

jhonyy9 (jhonyy9):

xsqrt8x =2xsqrt2x 5sqrt18x =15sqrt2x 3sqrt2x +2xsqrt2x -15sqrt2x =2xsqrt2x -12sqrt2x =(x-6)2sqrt2x

OpenStudy (anonymous):

ohh, so... 3*sqrt(2x) + x*sqrt(8x) - 5*sqrt(18x becomes: 3*sqrt(2x) + 2x*sqrt(2x) - 15*sqrt(2x). then you can factor out the sqrt(2x) to get: sqrt(2x) * (3 + 2x - 15) which results in: sqrt(2x) * (-12 + 2x)? thats another way I got it, would this be the way to write it?

jhonyy9 (jhonyy9):

so 3sqrt2x and -15sqrt2x because are the same terms you can assum these

OpenStudy (anonymous):

this is weird, I just went onto my Algebra solving program (it just gives you the answer without explaining) and it said that the answer is actually this: -12sqrt2x + 2xsqrt2x?

jhonyy9 (jhonyy9):

and after make the GCF - great common factor - what will be

jhonyy9 (jhonyy9):

yes is right but from this you can subtracting again the GCF what will be 2sqrt2x

OpenStudy (anonymous):

but how did it get that answer? neither of our solutions equaled -12sqrt2x + 2xsqrt2x?

jhonyy9 (jhonyy9):

check my calcules

OpenStudy (anonymous):

ohhh I see it now, sorry I missed that

OpenStudy (anonymous):

So this would be the whole answer: 3sqrt(2x)+ xsqrt(8x) - 5sqrt(18x) xsqrt8x =2xsqrt2x 5sqrt18x =15sqrt2x 3sqrt2x +2xsqrt2x -15sqrt2x =2xsqrt2x -12sqrt2x =(x-6)2sqrt2x

jhonyy9 (jhonyy9):

yes exactly

OpenStudy (anonymous):

But I had to break it up like this... 3sqrt(2x)+ xsqrt(8x) - 5sqrt(18x) xsqrt(8x) = 2xsqrt(2x) 5sqrt(18x) = 15sqrt(2x) 3sqrt(2x) + 2xsqrt(2x) -15sqrt(2x) = 2xsqrt(2x) -12sqrt(2x) =(x-6)2sqrt(2x)

OpenStudy (anonymous):

Okay cool, but what about the 1st part of the question, the one asking what Nakim missed... Does what I put sound good?

OpenStudy (anonymous):

I mean it could be this to: the answer given was sqrt(2x) * (-10x) looks like nakim combined 3 + 2x - 15 to get -10x. the 3 and the -15 could be combined because they're like terms. however, the 2x could not be combined because it was different from them. it included an x in the term which the others did not have. Therefore, since they are not like terms, the answer couldn't possibly be (2x)*(-10)?? would that be right?

jhonyy9 (jhonyy9):

so if you will check it there 3sqrt2x +2xsqrt2x not is equal 5xsqrt2x but Nakim have considered hence and Nakim have subtracted this 5xsqrt2x from 15sqrt2x and hence have got this result of -10xsqrt2x

OpenStudy (anonymous):

ahhh, thank you so much!! :)

jhonyy9 (jhonyy9):

yw was my pleasure good luck bye

OpenStudy (anonymous):

you really are a big help!

OpenStudy (anonymous):

bye